\begin{thebibliography}{2000}
%\magnification=\magstep1
%\parindent 0pt\parskip 4pt
\font \sc = cmcsc8
\parindent 0pt
%\parskip1pt

%\centerline{\bf{References}}
%\begin{list}
%\setlength\labelwidth{20pt}
%\setlength\leftmargin{0pt}
\bibitem[]{}

\vskip2pt\hskip-1.5truecm{\sc{Blachman, N. M.}} and {\sc{Few, L.}}

\bibitem[1963]{BlachmanFew} Multiple packing of spherical caps.
Mathematika {\bf{10}} (1963) 84--88. MR0155234, %(27 \#5170)
DOI~10.1112/S0025579300003405

\vskip2pt\hskip-1.5truecm{\sc{Blinovsky, V.}}

\bibitem[1999]{Blinovsky} Multiple packing of the Euclidean sphere.
IEEE Trans. Inform. Theory {\bf{45}} (1999) no. 4, 1334--1337. MR1686277,
DOI: 10.1109/18.761296

\vskip2pt\hskip-1.5truecm{\sc{Blinovsky, V.}} and {\sc{Litsyn, S.}}

\bibitem[2011]{BlinovskyLitsyn} New asymptotic bounds on the size of multiple packings of the Euclidean sphere.
Discrete Comput. Geom. {\bf{46}} (2011) no. 4, 626--635. MR2846172,
DOI~10.1007/s00454-011-9336-x

\vskip2pt\hskip-1.5truecm{\sc{Blundon, W. J.}}

\bibitem[1957]{Blundon57} Multiple covering of the plane by circles.
Mathematika {\bf{4}} (1957) 7--16. MR0090823, %(19,877b)
DOI~10.1112/S0025579300001042

\bibitem[1963]{Blundon63} Multiple packing of circles in the plane.
J. London Math. Soc. {\bf{38}} (1963) 176--182. MR0149374, %(26 \#6864)
DOI~10.1112/jlms/s1-38.1.176

\vskip2pt\hskip-1.5truecm{\sc{Bolle, U.}}

\bibitem[1976]{Bolle76} Mehrfache Kreisanordnungen in der euklidischen Ebene.
Thesis, Universit\"at Dortmund, 1976.

\bibitem[1979]{Bolle79} Dichteabsch\"atzungen f\"ur mehrfache gitterf\"ormige Kugelanordnungen im $R^m$.
Studia Sci. Math. Hungar. {\bf{14}} (1979) no. 1-3, 51--68. MR0645513

\bibitem[1982]{Bolle82} Dichteabsch\"atzungen f\"ur mehrfache gitterf\"ormige Kugelanordnungen im $R^m$ II.
Studia Sci. Math. Hungar. {\bf{17}} (1982) no. 1--4, 429--444. MR0761558

\bibitem[1984]{Bolle84} \"Uber die Dichte mehrfacher gitterf\"ormiger Kreisanordnungen in der Ebene.
Studia Sci. Math. Hungar. {\bf{19}} (1984) no. 2-4, 275--284. MR0874495

\bibitem[1989]{Bolle89} On the density of multiple packings and coverings of convex discs.
Studia Sci. Math. Hungar. {\bf{24}} (1989) no. 2--3, 119--126. MR1051142

\bibitem[1994]{Bolle94} On multiple tiles in $E^2$.
Intuitive geometry (Szeged, 1991), 39--43, Colloq. Math. Soc. J\'anos Bolyai, 63, North-Holland, Amsterdam, 1994. MR1383609

\vskip2pt\hskip-1.5truecm{\sc{Chan, Swee Hong}}

\bibitem[2015]{Chan} Quasi-periodic tiling with multiplicity: A lattice enumeration approach.
Discrete Comput. Geom. {\bf{54}} (2015) 647--662. MR3392970,
DOI~10.1007/s00454-015-9713-y

\vskip2pt\hskip-1.5truecm{\sc{Cohn, M. J.}}

\bibitem[1976]{CohnMJ} Multiple lattice coverings of space.
Proc. London Math. Soc. (3) {\bf{32}} (1976) 117--132. MR0389773,
DOI~10.1112/plms/s3-32.1.117

\vskip2pt\hskip-1.5truecm{\sc{Danzer, L.}}

\bibitem[1960]{Danzer60} Drei Beispiele zu Lagerungsproblemen.
Arch. Math. (Basel)  {\bf{11}} (1960) no. 1, 159--165. MR1552702,
DOI~10.1007/BF01236925

\vskip2pt\hskip-1.5truecm{\sc{Dumir, V. C.}} and Hans-Gill, R. J.

\bibitem[1972a]{DumirHans-Gill72a} Lattice double coverings in the plane.
Indian J. Pure Appl. Math. {\bf{3}} (1972) no. 3, 466--480. MR0308057 %(46 \#7172)

\bibitem[1972b]{DumirHans-Gill72b} Lattice double packings in the plane.
Indian J. Pure Appl. Math.  {\bf{3}} (1972) no. 3, 481--487. MR0308058 %(46 \#7173)

\vskip2pt\hskip-1.5truecm{\sc{Fejes T\'oth, G.}}

\bibitem[1976a]{FTG76a} Multiple packing and covering of the plane with circles.
Acta Math. Acad. Sci. Hungar.  {\bf{27}}  (1976) no. 1-2, 135--140. MR0417930, %(54 \#5977)
DOI~10.1007/BF01896768

\bibitem[1979]{FTG79} Multiple packing and covering of spheres.
Acta Math. Acad. Sci. Hungar. {\bf{34}} (1979) no. 1-2, 165--176. MR0546731, %(80k:52004)
DOI~10.1007/BF01902605

\bibitem[1984a]{FTG84a} Multiple lattice packings of symmetric convex domains in the plane.
J. London Math. Soc. (2) {\bf{29}} (1984)  no. 3, 556--561. MR0754941, %(86b:52009)
DOI~10.1112/jlms/s2-29.3.556

\vskip2pt\hskip-1.5truecm{\sc{Fejes T\'oth, L.}}

\bibitem[1966a]{FTL66a} Mehrfache Kreisunterdeckungen und Kreis\"uberdeckungen auf der Kugel.
Elem. Math. {\bf{21}} (1966) 34--35. MR0192414 %(33 \#639)

\vskip2pt\hskip-1.5truecm{\sc{Few, L.}}

\bibitem[1964]{Few64} Multiple packing of spheres.
J. London Math. Soc. {\bf{39}} (1964) 51--54. MR0161228, %(28 \#4436)
DOI~10.1112/jlms/s1-39.1.51

%\bibitem[1967a]{} Multiple packing of spheres: A survey. 1967 Proc. Colloquium on Convexity (Copenhagen, 1965) pp. 88--93 Kobenhavns Univ. Mat.
%Inst., Copenhagen. MR0215193 %(35 \#6036)

\bibitem[1967]{Few67} Double covering with spheres.
Mathematika {\bf{14}} (1967) 207--214. MR0228443, %(37 \#4023)
DOI~10.1112/S002557930000382X

\bibitem[1968]{Few68} Double packing of spheres: A new upper bound.
Mathematika {\bf{15}} (1968) 88--92. MR0234359, %(38 \#2676)
DOI~10.1112/S0025579300002424

\bibitem[1971]{Few71} The multiplicity of partial coverings of space.
Acta Arith. {\bf{18}} (1971) no. 1, 213--219. MR0290253, %(44 #7437)
DOI~10.4064/aa-18-1-213-219

\vskip2pt\hskip-1.5truecm{\sc{Few, L.}} and {\sc{Kanagasabapathy, P.}}

\bibitem[1969]{FewKanagasabapathy} The double packing of spheres.
J. London Math. Soc. {\bf{44}} (1969) 141--146. MR0243430, %(39 \#4752)
DOI~10.1112/jlms/s1-44.1.141

\vskip2pt\hskip-1.5truecm{\sc{Florian, A.}}

\bibitem[1978a]{Florian78a} Mehrfache Packung konvexer K\"orper.
\"Osterreich. Akad. Wiss. Math.-Natur. Kl. Sitzungsber. II {\bf{186}} (1978) no. 8--10,
373--384. MR0510352
    
\vskip2pt\hskip-1.5truecm{\sc{Frankl, N.; Nagy, J.}} and {\sc{Nasz\'odi, M.}}
    
\bibitem[2018]{FranklNagyNaszodi} Coverings: Variations on a result of
Rogers and on the Epsilon-net theorem of Haussler and Welzl.
Discrete Math. {\bf{341}} (2018) no. 3, 863--874. MR3754400,
DOI~10.1016/j.disc.2017.11.017

\vskip2pt\hskip-1.5truecm{\sc{Galiev, Sh. I.}}

\bibitem[1996]{Galiev} Multiple packings and coverings of a sphere. (Russian)
Diskret. Mat. {\bf{8}} (1996) no. 3, 148--160;
translation in Discrete Math. Appl. {\bf{6}} (1996) no. 4, 413--426. MR1422355,
DOI~10.4213/dm531
   
\vskip2pt\hskip-1.5truecm{\sc{Gravin, N.}}; {\sc{Kolountzakis, M.; Robins, S.}} and {\sc{Shiryaev, D.}}

\bibitem[2013]{GravinKolountzakisRobinsShiryaev} Structure results for multiple tilings in $3D$.
Discrete Comput. Geom. {\bf{50}} (2013) no. 4, 1033--1050. MR3138144,
DOI~10.1007/s00454-013-9548-3

\vskip2pt\hskip-1.5truecm{\sc{Gravin, N.}}; {\sc{Robins, S.}} and {\sc{Shiryaev, D.}}

\bibitem[2012]{GravinRobinsShiryaev} Translational tilings by a polytope, with multiplicity.
Combinatorica {\bf{32}} (2012) no. 6, 629--649. MR3063154,
DOI~10.1007/s00493-012-2860-3

\vskip2pt\hskip-1.5truecm{\sc{Groemer, H.}}

\bibitem[1978]{Groemer78}  On multiple space subdivisions by zonotopes.
Monatsh. Math. {\bf{86}} (1978/79) no. 3, 185--188. MR0517024,
DOI~10.1007/BF01659718
    
\bibitem[1986a]{Groemer86a} Multiple packings and coverings.
Studia Sci. Math. Hungar. {\bf{21}} (1986), no. 1-2, 189--200. MR0898856

\vskip2pt\hskip-1.5truecm{\sc{Haas, A.}}

\bibitem[1976]{Haas} Die d\"unnste siebenfache gitterf\"ormige \"Uberdeckung der Ebene durch kongruente Kreise.
Dissertation, Universit\"at Wien, 1976.

\vskip2pt\hskip-1.5truecm{\sc{Heppes, A.}}

\bibitem[1955]{Heppes55} \"Uber mehrfache Kreislagerungen.
Elem. Math. {\bf{10}} (1955) 125--127. MR0074015 %(17,523e)

\bibitem[1959]{Heppes59} Mehrfache gitterf\"ormige Kreislagerungen in der Ebene.
Acta Math. Acad. Sci. Hungar. {\bf{10}} (1959) 141--148. MR0105066, %(21\#3812)
DOI~10.1007/BF02063295

\vskip2pt\hskip-1.5truecm{\sc{Kolountzakis, M.}}

\bibitem[2000]{Kolountzakis} On the structure of multiple translational tilings by polygonal regions.
Discrete Comput. Geom. {\bf{23}} (2000) no. 4, 537--553. MR1753701,
DOI~10.1007/s004540010014

\bibitem[2021]{Kolountzakis21} Deciding multiple tiling by polygons in polynomial time.
Period.\ Math.\ Hungar. {\bf{83}} (2021) no. 1, 32--38. MR4260146,
DOI~10.1007/s10998-020-00361-y

\vskip2pt\hskip-1.5truecm{\sc{Lev, N.}} and {\sc{Liu, B.}}

\bibitem[2000]{LevLiu} Multi-tiling and equidecomposability of polytopes by lattice translates.
Bull. London Math. Soc. {\bf{51}} (2019) no. 6, 1079--1098. MR4041013,
DOI~10.1112/blms.12297

\vskip2pt\hskip-1.5truecm{\sc{Linhart, J.}}

\bibitem[1983]{Linhart83} Eine Methode zur Bestimmung der Dichte einer dichtesten gitterf\"ormigen $k$-fachen Kreispackung.
Arbeitsber. Math. Inst. Univ. Salzburg, 1983.

\vskip2pt\hskip-1.5truecm{\sc{Liu, B.,}}

\bibitem[2021]{Liu} Periodic structure of translational multi-tilings in the plane.
Amer. J. of Math.
{\bf{143}} (2021) no. 6, 1841--1862.
DOI~10.1353/ajm.2021.0047

%arXiv:1809.03440v2 [math.CO] 22 Oct 2019.


\vskip2pt\hskip-1.5truecm{\sc{Minkowski, H.}}

\bibitem[1897]{Minkowski97} Allgemeine Lehrs\"atze \"uber die konvexen Polyeder.
Nachr. Kgl. Ges. Wiss. G\"ottingen Math.-phys. Kl. 1897, 198--219.
Ges. Abh. Leipzig: Teubner, 1911. Bd. 2. S. 109--127.
Ausgew\"ahlte Arbeiten zur Zahlentheorie und zur Geometrie 121--139.
Teubner-Archiv zur Mathematik, vol 12. Springer, Vienna.
DOI~10.1007/978-3-7091-9536-9\_5

\vskip2pt\hskip-1.5truecm{\sc{Nasz\'odi, M.}} and {\sc{Polyanskii, A.}}

\bibitem[2018]{NaszodiPolyanskii} Approximating set multi-covers.
European J. Combin. {\bf{67}} (2018) 174--180. MR3707225,
DOI~10.1016/j.ejc.2017.08.001

\vskip2pt\hskip-1.5truecm{\sc{Pach, J.}}

\bibitem[1985]{Pach85} A note on plane covering.
Diskrete Geometrie, 3. Kolloq., Salzburg 1985, 207-216 (1985). Zbl 0573.52023

\bibitem[1986]{Pach86} Covering the plane with convex polygons.
Discrete Comput. Geom. {\bf{1}} (1986) no. 1, 73--81. MR0824109, %(87c:52024)
DOI~10.1007/BF02187684

\vskip2pt\hskip-1.5truecm{\sc{Pach, J.}} and {\sc{P\'alv\"olgyi, D.}}

\bibitem[2016]{PachPalvolgyi} Unsplittable coverings in the plane.
Advances in Mathematics {\bf{302}} (2016) 433--457. MR3545936,
DOI~10.1016/j.aim.2016.07.011

\vskip2pt\hskip-1.5truecm{\sc{Pach, J.; P\'alv\"olgyi, D.}} and {\sc{T\'oth, G.}}

\bibitem[2013]{PachPalvolgyiToth} Survey on decomposition of multiple coverings.
Geometry---Intuitive, Discrete, and Convex 219--257, Bolyai Society Mathematical
Studies vol. 24, J\'anos Bolyai Math. Soc., Budapest, 2013. MR3204561,
DOI~10.1007/978-3-642-41498-5\_9

\vskip2pt\hskip-1.5truecm{\sc{P\'alv\"olgyi, D.}}

\bibitem[2013]{Palvolgyi} Indecomposable coverings with unit discs.
arXiv:1310.6900v1 [math.MG] 25 Oct 2013
    
\vskip2pt\hskip-1.5truecm{\sc{P\'alv\"olgyi, D.}} and {\sc{T\'oth, G.}}

\bibitem[2010]{PalvolgyiToth} Convex polygons are cover-decomposable.
Discrete Comput. Geom. {\bf{43}} (2010) no. 3, 483--496. MR2587832,
DOI~10.1007/s00454-009-9133-y

\vskip2pt\hskip-1.5truecm{\sc{Purdy, G. B.}}

\bibitem[1973]{Purdy} The lattice triple packing of spheres in Euclidean space.
Trans. Amer. Math. Soc. {\bf{181}} (1973) 457--470. MR0377706,
DOI~10.1090/S0002-9947-1973-0377706-4

\vskip2pt\hskip-1.5truecm{\sc{Schmidt, W. M.}}

\bibitem[1961]{Schmidt61} Zur Lagerung kongruenter K\"orper im Raum.
Monatsh. Math. {\bf{65}} (1961) 154--158. MR0126215
DOI 10.1007/BF01307024

\vskip2pt\hskip-1.5truecm{\sc{Sriamorn, K.}}

\bibitem[2014]{Sriamorn14} Multiple lattice packings and coverings of the plane with triangles.
arXiv:1412.5096v2, 2014.

\bibitem[2016]{Sriamorn16} On the multiple packing densities of triangles.
Discrete Comput. Geom. {\bf{55}} (2016) no. 1, 228--242. MR3439266,
DOI~10.1007/s00454-015-9748-0

\vskip2pt\hskip-1.5truecm{\sc{Sriamorn, K.}} and {\sc{Wetayawanich, A.}}

\bibitem[2015]{SriamornWetayawanich15} On the multiple covering densities of triangles.
Discrete Comput. Geom. {\bf{54}} (2015) no. 3, 717--727. MR3392975,
DOI~10.1007/s00454-015-9711-0

\vskip2pt\hskip-1.5truecm{\sc{Subak, H.}}

\bibitem[1960]{Subak} Mehrfache gitterf\"ormige \"Uberdeckungen der Ebene durch Kreise.
Dissertation, Universit\"at Wien, 1960.
    
\vskip2pt\hskip-1.5truecm{\sc{Tardos, G.}} and {\sc{T\'oth, G.}}

\bibitem[2007]{TardosToth} Multiple coverings of the plane with triangles.
Discrete Comput. Geom. {\bf{38}} (2007) no. 2, 443--450. MR2343315,
DOI~10.1007/s00454-007-1345-4

\vskip2pt\hskip-1.5truecm{\sc{Temesv\'ari, \'A. H.}}

\bibitem[1984a]{Temesvari84a} Die d\"unnste gitterf\"ormige $5$-fache Kreis\-\"uberdeckung der Ebene.
Studia Sci. Math. Hungar. {\bf{19}}  (1984) no. 2-4, 285--298. MR0874496 %(88k:52024)

\bibitem[1984b]{Temesvari84b} Die d\"unnste 2-fache doppelgitterf\"ormige Kreis\"uberdeckung der Euklidischen Ebene.
Studia Sci. Math. Hungar. {\bf{19}} (1984) no. 1, 109--116. MR787792 %(86j:52023)

\bibitem[1988]{Temesvari88} Eine Methode zur Bestimmung der d\"unnsten gitterf\"ormigen $k$-fachen Kreis\-\"uber\-deckungen.
Studia Sci. Math. Hungar. {\bf{23}} (1988)  no. 1-2, 23--35. MR0962429 %(89j:52014)

\bibitem[1992a]{Temesvari92a} Die d\"unnste gitterf\"ormige 6-fache Kreis\"uberdeckung.
Berzsenyi D\'aniel Tan\'ark. F\"oisk. Tud. K\"ozl., Termtud. {\bf{8}} (1992) 93-112. Zbl 0806.52015

\bibitem[1992b]{Temesvari92b} Die d\"unnste gitterf\"ormige 7-fache Kreis\"uberdeckung.
Berzsenyi D\'aniel Tan\'ark. F\"oisk. Tud. K\"ozl., Termtud. {\bf{8}} (1992) 113-125. Zbl 0806.52016

\bibitem[1994a]{Temesvari94a} Die d\"unnste $8$-fache gitterf\"ormige Kreis\-\"uberdeckung der Ebene.
Studia Sci. Math. Hungar. {\bf{29}} (1994) no. 3-4, 323--340. MR1304886 %(95i:52022)

\bibitem[1994b]{Temesvari94b} Die dichteste gitterf\"ormige $9$-fache Kreispackung.
Rad Hrvatske Akad. Znan. Umjet. No. 467  (1994) 95--110. MR1362504 %(96m:52030)

\bibitem[1994c]{Temesvari94c} \"Uber die d\"unnste doppelgitterf\"ormige
2-fache \"Uberdeckung mit einem zentralsymmetrischen konvexen Bereich.
%[On the thinnest 2-fold double-lattice covering by a centro-symmetric
%convex-domain]
Festschrift on the occasion of the 65th birthday of Otto Kr\"otenheerdt.
Beitr\"age Algebra Geom. {\bf{35}} (1994) no. 1, 45--54. MR1287196 %(95i:52016)
    
\vskip2pt\hskip-1.5truecm{\sc{Temesv\'ari, \'A. H.}}; {\sc{Horv\'ath, J.}} and {\sc{Yakovlev, N. N.}}

\bibitem[1987]{TemesvariHorvathYakovlev} A method for finding the densest lattice $k$-fold packing of circles. (Russian)
Mat. Zametki  {\bf{41}}  (1987)  no. 5, 625--636, 764.
English translation: Math. Notes {\bf{41}} (1987) no. 5-6, 349--355.  MR0898123, %(88i:11041)
DOI~10.1007/BF01159857

\vskip2pt\hskip-1.5truecm{\sc{Temesv\'ari \'A. H.}} and {\sc{V\'egh, A.}}

\bibitem[1998]{TemesvariVegh} Die dichteste gitterf\"ormige $10$-fache Kreispackung.
Berzsenyi D\'aniel Tan\'ark. F\"oisk. Tud. K\"ozl., Termtud. {\bf{6}} (1998) 3--20.

\vskip2pt\hskip-1.5truecm{\sc{Yakovlev, N. N.}}

\bibitem[1983]{Yakovlev83} The densest lattice $8$-packing on a plane. (Russian)
Vestnik Moskov. Univ. Ser. I Mat. Mekh. (1983) no. 5, 8--16.
English translation in Moscow Univ. Math. Bull. {\bf{38}} (1983) no. 5, 7--16. MR0722442 %(85m:52010)

\bibitem[1984]{Yakovlev84} A method of finding the densest lattice $k$-packing on a plane. (Russian)
Functional analysis and its applications in mechanics and probability theory (Moscow, 1983) 170--171,
Moskov. Gos. Univ., Moscow, 1984. MR0845122 %(87h:52028)

\vskip2pt\hskip-1.5truecm{\sc{Yang, Q.}}

\bibitem[2019]{Yang} On multiple translative tiling in the plane.
arXiv:1809.04272v4 [math.MG] 30 Sep 2019.

\vskip2pt\hskip-1.5truecm{\sc{Yang, Q.}} and {\sc{Zong, C.,}}

\bibitem[2019]{YangZong19}
Multiple lattice tilings in Euclidean spaces.
Canad.\ Math.\ Bull. {\bf{62}} (2019) no. 4, 923--929. MR4028498,
DOI~10.4153/S0008439518000103

\bibitem[2021]{YangZong21}
Characterization of the two-dimensional five-fold translative tiles.
Bull.\ Soc.\ Math.\ France {\bf{149}} (2021) no. 1, 119--153.



\end{thebibliography}
\end{document}

\end{document}

\documentclass{amsart}
\pdfoutput=1
\usepackage{caption}
\usepackage{subcaption}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{wrapfig}
\begin{document}

\author{G\'abor Fejes T\'oth}
\address{Alfr\'ed R\'enyi Institute of Mathematics,
Re\'altanoda u. 13-15., H-1053, Budapest, Hungary}
\email{gfejes@renyi.hu}

\title{Multiple arrangements}
\thanks{The English translation of the book ``Lagerungen in der Ebene,
auf der Kugel und im Raum" by L\'aszl\'o Fejes T\'oth will be
published by Springer in the book series Grundlehren der
mathematischen Wissenschaften under the title
``Lagerungen---Arrangements in the Plane, on the Sphere and
in Space". Besides detailed notes to the original text the
English edition contains eight self-contained new chapters
surveying topics related to the subject of the book but not
contained in it. This is a preprint of one of the new chapters.}

\begin{abstract}
This paper surveys the theory of multiple packings and coverings. The study
of multiple arrangements started in the 60s of the last century, and it
was restricted mostly to lattice arrangements on the plane or of general
arrangements of balls. We emphasize two new topics which were intensively
investigated recently: decomposition of multiple coverings into simple
coverings and characterization of multiple tilings.
\end{abstract}

\maketitle

We say that a family of sets is a {\it $k$-fold packing} if every point of the space
belongs to the interior of at most $k$ sets. Quite analogously, we say that a family
of sets forms a {\it $k$-fold covering} if every point of the space belongs to
the closure of at least $k$ sets. Let $\delta^k(K)$ and $\vartheta^k(K)$, respectively,
denote the densities of the densest $k$-fold packing and the thinnest $k$-fold
covering of the space with congruent copies of the convex body $K$. Similarly we use
the notation $\delta^k_T(K)$, $\vartheta^k_T(K)$, $\delta^k_L(K)$ and
$\vartheta^k_L(K)$ for the optimum densities of the corresponding $k$-fold translative
and lattice arrangements (compare the corresponding definitions of $\delta$, $\delta_T$,
$\delta_L$, $\vartheta$, $\vartheta_T$ and $\vartheta_L$ above).

\section{Multiple arrangements on the plane}

The literature on multiple packing and covering is relatively extensive,
and it mostly deals with arrangements of congruent copies of the circular disk
$B^2$. The values of $\delta^k(B^2)$ and
$\vartheta^k(B^2)$ are not known for any $k>1$. We know that
$\delta^2(B^2)>2\delta^1(B^2)$ and $\vartheta^2(B^2)<2\vartheta^1(B^2)$ as was shown by
{\sc{Heppes}}~\cite{Heppes55} and {\sc{Danzer}}~\cite{Danzer60}, respectively.
{\sc{G.~Fejes T\'oth}}~\cite{FTG76a} established the bounds
$$
\delta^k(B^2)\le k\,\frac{\pi}{6}\cot\frac{\pi}{6k}\quad{\rm{and}}\quad
\vartheta^k(B^2)\ge k\,\frac{\pi}{3}\csc\frac{\pi}{3k}.
$$
Observe that $\frac{\pi}{6}\cot\frac{\pi}{6k}$ and $\frac{\pi}{3}\csc\frac{\pi}{3k}$
are equal to the density of a disk with respect to the circumscribed and inscribed regular $6k$-gon.
For $k=1$ these inequalities are sharp as they coincide with Thue's and Kershner's theorems.

{\sc{Bolle}}~\cite{Bolle76} proved that there are positive constants $c_i$ such
that
$$k-c_1k^\frac{2}{5}\le\delta_L^k(B^2)\le{k}-c_2k^\frac{1}{4}$$
and
$$k+c_3k^\frac{1}{4}\le\vartheta_L^k(B^2)\le{k}+c_4k^\frac{2}{5}$$
and showed in~\cite{Bolle84} that the exponent $\frac{1}{4}$ is best possible in
these inequalities. In~\cite{Bolle89} he proved that for convex disks $K$
with piecewise twice differentiable boundary there are positive constants
$c(K)$ and $C(K)$ such that
$$\delta_L^k(K)\ge{k}-c(K)k^\frac{2}{5}$$
and
$$\vartheta_L^k(K)\le{k}+C(K)k^\frac{2}{5}.$$
Moreover, for a polygon $P$ the stronger inequalities
$$\delta_L^k(P)\ge{k}-c(P)k^\frac{1}{3}$$
and
$$\vartheta_L^k(P)\le{k}+C(P)k^\frac{1}{r3}$$
hold.

The exact values of $\delta^k_L(B^2)$ have been found for $k\le10$
(see {\sc{Heppes}}~\cite{Heppes59}, {\sc{Blundon}}~\cite{Blundon63},
{\sc{Bolle}}~\cite{Bolle76}, {\sc{Yakovlev}}~\cite{Yakovlev83}, {\sc{Temesv\'ari}}
~\cite{Temesvari94b} and {\sc{Te\-mes\-v\'ari}} and {\sc{V\'egh}}~\cite{TemesvariVegh}). The values of
$\vartheta^k_L(B^2)$ are known for $k\le8$ (see {\sc{Blundon}}~\cite{Blundon57}, {\sc{Haas}}~\cite{Haas},
{\sc{Subak}}~\cite{Subak} and {\sc{Temesv\'ari}}~\cite{Temesvari84a,Temesvari92a,Temesvari92b,Temesvari92a,Temesvari94a}.
{\sc{Linhart}}~\cite{Linhart83} described an algorithmic
approach for approximating the values of $\delta^k_L(B^2)$ with arbitrarily high
accuracy. Elaborating on results by {\sc{Yakovlev}}~\cite{Yakovlev84} {\sc{Temesv\'ari,
Horv\'ath}} and {\sc{Yakovlev}}~\cite{TemesvariHorvathYakovlev} described a method for finding the
densest $k$-fold lattice packing with circles. They reduced this task to a
finite number of optimization problems, each over an explicitly given compact
domain. A similar method for the thinnest $k$-fold lattice covering with circles
was given by {\sc{Temesv\'ari}}~\cite{Temesvari88}.

For a triangle $T$, {\sc{Sriamorn}}~\cite{Sriamorn14} determined $\delta_L^k(T)$ and
$\vartheta_L^k(T)$ for all $k$. Moreover, {\sc{Sriamorn}}~\cite{Sriamorn16}
showed that
$$\delta_T^k(T)=\delta_L^k(T)=\frac{2k^2}{2k+1}$$
and {\sc{Sriamorn}} and {\sc{Wetayawanich}}~\cite{SriamornWetayawanich15} showed that
$$\vartheta_T^k(T)=\vartheta_L^k(T)=\frac{2k+1}{2}$$
for all $k$.

It is worth mentioning that $\delta^k_L(B^2)=k\delta^1_L(B^2)$ for $k=2,\ 3$ and
$4$, and also $\vartheta^2_L(B^2)=2\vartheta^1_L(B^2)$. These equalities for the
very same multiplicities have been extended to an arbitrary centrally symmetric
convex disk in place of the circle by {\sc{Dumir}} and {\sc{Hans-Gill}}
\cite{DumirHans-Gill72a,DumirHans-Gill72b} and {\sc{G.~Fejes T\'oth}} \cite{FTG84a}.
The equality $\vartheta^2_L(B^2)=2\vartheta^1_L(B^2)$ was further generalized by
{\sc{Temesv\'ari}}, who proved in \cite{Temesvari84b} that the density of a 2-periodic
double covering by circles is at most $2\vartheta^1_L(B^2)$, and in \cite{Temesvari94c}
proved the analogous result for centrally symmetric convex disks. Recall from XX that
an ${m}$-periodic arrangement is the union of $m$ translates of a lattice arrangement.

\section{Decomposition of multiple arrangements}

The equalities $\delta^3_L(K)=3\delta^1_L(K)$ and $\delta^4_L(K)=4\delta^1_L(K)$
for centrally symmetric disks $K$ were derived by noticing that every $3$-fold lattice
packing by such a disk is the union of three simple ($1$-fold) packings, and every such
$4$-fold packing is the union of two $2$-fold packings. This observation belongs to
the topic concerning decompositions of multiple arrangements into simple ones,
problems and results that focus on the combinatorial structure of such arrangements.
Research in this direction was initiated by {\sc{Pach}}~\cite{Pach85}. He proved,
among other things, that every $2$-fold packing with positively homothetic copies of
a convex disk can be decomposed into four (simple) packings. For coverings, he made
the conjecture that for every convex disk $K$ there exists a minimal natural
number $m(K)$ such that every $m(K)$-fold covering of the plane by translates
of $K$ can be decomposed into two coverings. In \cite{Pach86} he proved this
conjecture for centrally symmetric polygons. New interest arose in the topic
after {\sc{Tardos}} and {\sc{T\'oth}}~\cite{TardosToth} proved the conjecture
for triangles. Soon after, {\sc{P\'alv\"olgyi}} and {\sc{T\'oth}}
~\cite{PalvolgyiToth} proved the conjecture for every convex polygon $P$.
Unfortunately, the number $m(P)$ increases with the number of sides of $P$,
thus the attempt to extend the result to all convex disks through polygonal
approximation fails. Still, it came as a surprise when {\sc{P\'alv\"olgyi}}
\cite{Palvolgyi} (see also {\sc{Pach}} and {\sc{P\'alv\"olgyi}}
\cite{PachPalvolgyi}) disproved Pach's conjecture by showing that it does
not hold for the circle. For subsequent developments on decomposition of
multiple arrangements we refer the reader to the survey article of
{\sc{Pach, P\'alv\"olgyi}} and {\sc{T\'oth}} \cite{PachPalvolgyiToth}.

\section{Multiple arrangements in space}

The densest 2-fold lattice packing and the thinnest 2-fold lattice covering of
balls in three dimensions were determined by {\sc{Few}} and {\sc{Kanagasabapathy}}
\cite{FewKanagasabapathy} and {\sc{Few}} \cite{Few67},
respectively. {\sc{Purdy}} \cite{Purdy} constructed a threefold lattice packing
of balls which he conjectured to be of maximum density. He supported the
conjecture by proving that it provides a local maximum of the density among
threefold lattice packings of balls.

Adapting Blichfeldt's idea, {\sc{Few}} \cite{Few64} gave the
following upper bound for the $k$-fold packing density of the $n$-dimensional
ball:
$$
\delta^k(B^n)\le (1+n^{-1})[(n+1)^k  -1][k/(k+1)]^{n/2}.
$$
This is better than the trivial bound $k$ only for large values of $n$ compared
to $k$. By a further elaboration on the same idea for $k=2$, {\sc Few}
\cite{Few68} obtained the stronger inequality
$$
\delta^2(B^n)\le\frac{4}{3}(n+2)\left(\frac{2}{3}\right)^{n/2}.
$$

{\sc G.~Fejes T\'{o}th} \cite{FTG79} gave a non-trivial upper bound for
$\delta^k(B^n)$, as well as a non-trivial lower bound for $\vartheta^k(B^n)$
for every $n$ and $k$.

For multiple lattice arrangements of balls, {\sc{Bolle}} \cite{Bolle79,Bolle82}
established sharper estimates. He proved that there are positive constants
$c_n$ and $C_n$ such that
$$
\frac{\delta_L^k(B^n)}{k}\le1-c_nk^{n+\frac{1}{n}}\quad{\hbox{and}}\quad
\frac{\vartheta_L^k(B^n)}{k}\ge1+c_nk^\frac{n+1}{2n}
$$
when $n\not\equiv1$ (mod 4)
and
$$
\frac{\delta_L^k(B^n)}{k}\le1-c_nk^\frac{n+3}{2n}\quad{\hbox{and}}\quad
\frac{\vartheta_L^k(B^n)}{k}\ge1+c_nk^\frac{n+3}{2n}
$$
when $n\equiv1$ (mod 4).

Extending the result of {\sc{Schmidt}} \cite{Schmidt61} to multiple
arrangements, {\sc{Florian}} \cite{Florian78a} proved that
$\delta^k(K)<k$ and $\vartheta^k(K)>k$ for every smooth
convex body $K$ without establishing a concrete bound.

By a Blichfeldt-type argument {\sc{Few}} \cite{Few64} proved
$$
\delta^k(B^n)\ge\delta(B^n)\left(\frac{2k}{k+1}\right)^{n/2}.
$$
In \cite{Few71} {\sc{Few}} studied the multiplicity of partial
coverings of space, and, as an application of a general theorem,
obtained a better lower bound for $\delta^k(B^n)$ for large values
of $k$ and $n$.

{\sc{Groemer}} \cite{Groemer86a} proved lower bounds for the $k$-fold
lattice packing density of a convex body $K$ involving the intrinsic
volumes of $K$. From his results follows the existence of positive
constants $c_n$ such that
$$\delta_L^k(K)\ge k-c_nk^{(n-1)/n}$$
for every convex body $K\in{E^n}$.

{\sc{Cohn}} \cite{CohnMJ} proved that
$$\vartheta_L^k(K)<[{(k+1)}^{1/n}+8n]^n=k(1+O(n^2k^{-1/n}))\qquad{\rm{as}}\ k\to\infty$$
for every $n$-dimensional convex body $K$.

The best known upper bound for $\vartheta^k(K)$ is due to {\sc{Nasz\'odi}}
and {\sc{Polyanskii}} \cite{NaszodiPolyanskii} who, improving slightly
on an earlier result by {\sc{Frankl, Nagy}} and {\sc{Nasz\'odi}}
\cite{FranklNagyNaszodi}, proved that
$$\vartheta^k(K)\le3.153(1+o(1))\max\{n\ln{n},k\}$$
for every convex body $K\in E^n$.

{\sc{Blachman}} and {\sc{Few}} \cite{BlachmanFew} gave bounds for the density
of multiple packings of spherical caps. Also {\sc L.~Fejes T\'oth} \cite{FTL66a},
{\sc{Galiev}} \cite{Galiev}, {\sc{Blinovsky}} \cite{Blinovsky} and {\sc{Blinovsky}}
and {\sc{Litsyn}} \cite{BlinovskyLitsyn} investigated multiple ball packings
in spherical spaces.

\section{Multiple tiling}

A system of bodies forms a {\it{$k$-fold tiling}} if each point of the space is
covered exactly $k$ times, except perhaps the boundary points of the bodies.
There are centrally symmetric polytopes that admit a translational
$k$-fold tiling, but no simple tiling. The simplest example is perhaps the
regular octagon of side-length 1, whose translates by the unit square
lattice form a 7-fold tiling.

{\sc{Bolle}} \cite{Bolle94} proved that a convex polygon that admits a
$k$-fold tiling of the plane by translations is centrally symmetric.
He also gave a characterization of those convex polygons that admit a
$k$-fold lattice-tiling. {\sc{Kolountzakis}} \cite{Kolountzakis21} gave an algorithm
which decides for a centrally symmetric convex polygon if it can tile the
plane by translations at some level. His algorithm runs in polynomial time
in the number of sides of the polygon. {\sc{Yang}} and {\sc{Zong}}
\cite{YangZong19} \cite{YangZong21} characterized those convex polygons that
admit two-, three-, four- or five-fold translational tiling. Only parallelograms
and centrally symmetric hexagons admit a two-, three- or four-fold translational
tiling. There are two more classes of polygons admitting five-fold tilings: the
affine images of a special octagon and of a decagon.

{\sc{Gravin, Robins}} and {\sc{Shiryaev}} \cite{GravinRobinsShiryaev}
proved that if translates of
a convex polytope form a $k$-fold tiling of $E^n$, then it is centrally
symmetric and its facets are centrally symmetric as well. This generalizes
a theorem of {\sc{Minkowski}} \cite{Minkowski97} concerning simple tilings.
For the three-dimensional case this means that only zonotopes admit a
translational $k$-fold tiling. For rational polytopes Gravin, Robins and
Shiryaev also proved the converse of their above mentioned theorem: Every
rational polytope in $E^n$ that is centrally symmetric and has centrally symmetric
facets admits a $k$-fold lattice tiling for some positive integer $k$.
For zonotopes, this was proved earlier by {\sc{Groemer}} \cite{Groemer78}.

A {\it{quasi-periodic}} set is a finite union of translated lattices, not
necessarily of the same lattice. {\sc{Kolountzakis}} \cite{Kolountzakis}
proved that a $k$-fold tiling by translates of a convex polygon other
than a parallelogram is quasi-periodic. {\sc{Gravin, Kolountzakis, Robins}}
and {\sc{Shiryaev}} \cite{GravinKolountzakisRobinsShiryaev} proved an
analogous theorem for the three-dimensional case: A $k$-fold tiling by
translates of a polytope that is not a two-flat zonotope is
quasi-periodic. A {\it{two-flat zonotope}} is the Minkowski sum of
two 2-dimensional symmetric polygons one of which may degenerate into
a single line segment.

{\sc{Gravin, Robins}} and {\sc{Shiryaev}} \cite{GravinRobinsShiryaev}
raised the problem whether the following generalization of the
Venkov-McMullen theorem holds: If translates of a polytope $P$ form a
$k$-fold tiling of $E^n$, then $P$ also admits an $m$-fold lattice
tiling for some, possibly different, multiplicity.  The
two-dimensional case of this conjecture was confirmed independently
by {\sc{Liu}} \cite{Liu} and {\sc{Yang}} \cite{Yang}. A further step in
the direction of proving the conjecture was made by {\sc{Chan}} \cite{Chan},
who proved it for certain quasi-periodic $k$-fold tilings. {\sc{Lev}}
and {\sc{Liu}} \cite{LevLiu} gave a characterization of those polytopes
in $E^n$ that tile with some multiplicity $k$ by translations along a
given lattice.

\small{
\bibliography{pack}}}
\mathrm{}